OpenStudy (anonymous):

Explain, in complete sentences, how you would factor 8x6 - 56x3 + 98.

OpenStudy (cwrw238):

first divide through by GCF = 4 = 2(4x6 - 28x3 + 49) now try factor expression in brackets note 4 = 2*2 and 49 = -7*-7 and middle term is -28 this is perfect square and = (2x^3 - 7)^2 factors are 2(2x^3 - 7)(2x^3 - 7)

OpenStudy (cwrw238):

* GCF = 2